OpenStudy (anonymous):

write 27^-2/3 in radical notation. Evaluate radical expression

OpenStudy (shamil98):

\[\huge \sqrt[n]{x^m} = x^\frac{ m }{ n }\]

OpenStudy (ranga):

Write 27 as 3^3 and simplify.

OpenStudy (shamil98):

\[\huge x^{-n} = \frac{ 1 }{ x^n }\]

OpenStudy (shamil98):

So using these formulas you know that this is : \[\huge \frac{ 1 }{ \sqrt[3]{27^2} }\]

OpenStudy (shamil98):

From there just simplify.

OpenStudy (anonymous):

I am horrible at this stuff. It looks like gibrish to me. so if i do 3^3 and that equals 27 its my answer?

OpenStudy (shamil98):

\[\huge \frac{ 1 }{ \sqrt[3]{27^2} } = \frac{ 1 }{ \sqrt[3]{729} }\] Just simplify it.

OpenStudy (shamil98):

What value multiplied by itself results in 729?..

OpenStudy (shamil98):

by itself three times*

OpenStudy (anonymous):

9

OpenStudy (shamil98):

Correct. So your resulting evaluated radical is \[\huge \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

that makes more sense now. thankyou